If $100 is borrowed and the interest after 6 months is $8, what is the annual interest rate for a simple interest loan?
Rate =

Answer by math_tutor2020(3817)


Answer: 16%

Work Shown:
i = P*r*t
8 = 100*r*(6/12)
8 = 100*r*(0.5)
8 = 50r
r = 8/50
r = 16/100
r = 0.16
Move the decimal point two spots to the right to arrive at an annual simple interest rate of 16%

Take note how t = 6/12 = 0.5 since t is the number of years.